# Evaluating Quantumness, Efficiency and Cost of Quantum Random Number Generators via Photon Statistics

### v5

@@ Abstract
From the user’s perspective, the fundamental parameters of a QRNG are quantumness, efficiency (random entropy and random number generation rate), and cost. Our analysis reveals that these parameters depend on three factors, namely, expected photon count per unit time, external reference cycle duration, and detection efficiency. A lower expected photon count enhances entropy but increases cost and decreases the generation rate. A shorter external reference cycle boosts entropy but must exceed a minimum threshold to minimize timing errors, with minor impacts on cost and rate. Lower detection efficiency enhances entropy and lowers cost but reduces the generation rate. Finally, to validate our results, we perform statistical tests like NIST, Dieharder, AIS-31, ENT etc. over the data simulated with different values of the above parameters. Our findings can empower manufacturers to customize QRNGs to meet user needs effectively.
@@ I Introduction
Random numbers have important roles in numerous applications, including computational methods like Monte Carlo simulations and programming RevModPhys.87.1067 ; MonteCarlo , testing principles of physics and chemistry PhysicsPhysiqueFizika.1.195 , as well as in the expansive realm of cryptography for creating cryptographic codes or concealing messages Bennett ; Crypto , extending to commercial uses such as slot machines and lottery games konstantinou2004electronic ; yamanaka2017enumeration . In cryptography, randomness is a key input to ensure security in communication. Classical pseudo-random number generators (PRNG) cannot guarantee communication security due to their deterministic generation process. On the contrary, the randomness of a true random number generator (TRNG) is rooted in some physical entropy source Schindler2009 ; Stipčević2014 , which may be classical or quantum.
The randomness of a classical true random number generator (CTRNG) depends on the unpredictability of the physical process that generates the entropy. This unpredictability arises from our limited knowledge of the system’s behavior Stipčević2014 ; mannalatha2023comprehensive . However, if future research resolves these limitations, the process could become predictable and would no longer function as a true random number generator. In contrast, a quantum true random number generator (QTRNG, or more commonly, QRNG) exploits the properties of quantum particles and the principles of quantum mechanics ma2016quantum to produce random numbers. According to quantum theory, these numbers are fundamentally random, making them inherently unpredictable. Thus it would remain unpredictable, irrespective of any technological advancement.
@@
Different technologies, including photon arrival time 10.1063/1.2720728 ; 10.1063/1.2961000 ; WayneRNG ; WayneQRNG ; 10.1063/1.3578456 ; 10.1063/1.4863224 ; 7270265 ; zhang2021simple , tunneling bernardo2017extracting ; 10.1063/5.0055955 , fluctuation gabriel2010generator ; PhysRevApplied.3.054004 ; Guo:19 ; ZQVR ; 10.1063/1.5078547 ; gehring2021homodyne , phase noise Qi:10 ; PhysRevE.81.051137 ; PhysRevA.81.063814 ; Xu:12 ; 10.1063/1.4922417 ; Yang:16 ; app10072431 , etc. have been proposed to produce QRNGs RevModPhys.89.015004 ; mannalatha2023comprehensive . Several manufacturers such as ID Quantique quantis , Qutools quRNG , QNuLabs TROPOS , QuintessenceLabs qStream , Mars Innovation quantoss , Quantum Computing Inc. uQRNG , Quside quside , QuantumCTek QCTek are producing QRNGs following these technologies. These QRNGs are commercially available as off-the-shelf black-box devices. As the photon arrival-time-based QRNGs are the most general and most of companies follow this approach, in this work we focus on the analysis of this type of QRNGs only.
@@
Since no black-box randomness-testing algorithm can distinguish between a PRNG and a QRNG, therefore the first step in validating the quantumness [^1] Sub-Poissonian distribution ($mean>variance$) of photon statistics (more details in Section II). of a random number generator (RNG) is to collect information regarding intermediate attenuator and detector parameters and photon statistics. Then it may be possible to devise a strategy to decide whether the device performs in the classical or quantum regime.
In most of the analyses Jones:23 ; Perina:24 ; Perina24 ; PhysRevA.109.043505 of the photon arrival-time-based QRNGs, the photon statistics have been discussed either by comparing the mean and variance of the sample distribution or using the negativity of the Mandel parameter Mandel:79 . For example, Alléaume et al. Alléaume_2004 discussed photon statistics of photon sources using the Mandel parameter given by (Mandel:79, , Eq. (11a)). They calculated values of the Mandel parameter from observed photon count timestamps and concluded that negative values of the Mandel parameter, which indicates sub-Poissonian statistics, signify non-classical photon statistics for photon sources.
Although the above strategy is analytically correct, a negative value of the Mandel parameter for experimental data may not always indicate sub-Poissonian statistics. For example, a finite number of photon count timestamps from coherent light that actually follows Poissonian distribution may result in a negative value of the Mandel parameter. This happens because the mean and the variance of a sample with a finite size may not always be the same (even under approximation) as the mean and the variance of the underlying distribution BERG2005865 . Therefore, calculating the exact value of any parameter from a sample of finite size is not enough to decide the underlying nature of the distribution.
Another approach to test the quantumness of light is to observe photon antibunching RevModPhys.54.1061 experimentally. However, although a stream of antibunched photons indicates the non-classical nature of light, the converse is not necessarily true. In other words, non-classical light can sometimes be observed as bunched photons PhysRevA.41.475 ; fox2006quantum . We explain this in more detail inB. This approach also consists of equality criteria similar to the above photon statistics case. Therefore this is also not enough to decide photon antibunching.
On the other hand, a single-photon source ensures the quantumness of a QRNG RevModPhys.54.1061 ; fox2006quantum . However, generating single photons are experimentally challenging task PhysRevLett.72.210 ; kim1999single . Various experimental methods PhysRevLett.83.2722 ; lounis2000single ; PhysRevLett.86.1502 ; santori2002indistinguishable ; 10.1063/1.1577828 ; PhysRevLett.89.067901 ; adam2024single ; wang2024multimode have been explored to produce single photons. Existing time-bin-based QRNG models utilize coherent photon sources in conjunction with attenuators to produce random numbers from exponential WayneQRNG ; WayneRNG ; 10.1063/1.3578456 or uniform 10.1063/1.4863224 distributions if the underlying optical devices perform ideally and the source generates perfectly single photon. But, in practice, this is not the case. Nie et al. 10.1063/1.4863224 shows that, for a photon arrival-time-based QRNG, the uniformity and hence the min-entropy depends on device parameters such as expected photon number within a fixed time interval as well as on the detection efficiency. Therefore, it is important to explore photon statistics in practical scenarios in more detail.
The QRNG output theoretically follows a perfectly uniform distribution. However, due to the imperfections of the underlying devices, the raw QRNG data may deviate from uniform randomness and therefore require classical post-processing. One may note that deterministic classical post-processing does not increase the randomness of the QRNG output, it only compresses the raw data with an aim to extracting the min-entropy present in it at the cost of the randomness generation rate.
@@ I.1 Motivation
Also, there are several schemes of device-independent QRNG (DI-QRNG) pironio2010random ; 10.1145/2885493 ; acin2016certified ; bierhorst2018experimentally ; liu2018device ; PhysRevLett.120.010503 ; 9153851 and certified randomness pironio2010random ; 6326426 ; Passaro_2015 ; acin2016certified ; Nguyen:18 ; 10.1093/oso/9780198788416.003.0008 ; PhysRevX.10.041048 ; zhang2021simple ; Leone:23 ; PhysRevLett.133.020802 ; PRXQuantum.5.020348 based on Bell violation, non-separability, etc. acin2016certified ; Alicki_2008 ; Alicki_2008a ; Martin_2023 . Though they do not rely on trusted device assumptions and do not require any post-processing, the required resources (that is, entangled qubits) are too costly. More recent works Mironowicz_2016 ; 10.1145/3564246.3585145 attempt locally certified randomness without entanglement, but they require random seeds and some kind of bootstrapping that may be considered analogous to classical post-processing.
On the other hand, the optical trusted device QRNG mannalatha2023comprehensive relies on the assumption that the devices used for extracting randomness and the source are well-characterized and no one can change any configuration. This means that no third party can access the QRNG or influence its results. In contrast, single-photon-based QRNGs can still be verified using photon-count statistics PhysRevA.41.475 ; fox2006quantum ; RevModPhys.68.127 ; qute.202100062 improving the security. A detailed comparison between trusted QRNG and DI-QRNG is given in A. In this work, we focus on the QRNG of the optical trusted device type along with a two-fold statistical method to verify its quantumness.
Side-channel attacks 8924447 ; 8942152 ; chowdhury2021physical can be performed in the post-processing step to gather information about the random number. Therefore, to reduce the compression amount of raw data and to prevent the data from side-channel attacks, the amount of post-processing should be minimal to ensure the maximum entropy in raw data with high speed and reduce the possibility of side-channel attacks on post-processing. To achieve this, the raw output number (before post-processing) generated by QRNG should be as close to uniform random as possible. Nie et al. 10.1063/1.4863224 provided a lower bound for the min-entropy of the raw data in the presence of device errors. Depending on this min-entropy, a suitable post-processing algorithm can be chosen to retain the entropy from the quantum source. As all the practical devices are prone to error, this raises an important question:
@@ I.2 Our Contributions
In this paper, we aim to evaluate the quantumness of photon arrival-time-based QRNG from the manufacturer’s perspective as answers to Q1 and Q2 and randomness, random number generation rate, and cost from the user’s perspective as the answer to Q3. A similar analysis can also be performed for other QRNG models. For the manufacturer, we first address the validation of quantum sources using photon counting statistics. Considering that the sample data for the number of photons within a fixed time is available, we suggest a two-fold statistical method to verify whether the source of the sample is quantum or not, overcoming the limitations of the existing approaches discussed above. The method consists of interval estimation and hypothesis testing, to ensure sub- or super-Poissonian distribution up to a desired confidence level. Several photon-number-resolving detectors have been proposed PhysRevA.71.033819 ; Divochiy2008 ; Kardyna2008 ; PhysRevA.99.043822 ; Natarajan2012 ; Schuck2013 ; Hadfield2009 ; 9844820 ; Provaznik:20 ; Marsili2009 and demonstrated 1263782 ; PhysRevA.67.061801 ; 10.1063/1.1596723 ; PhysRevA.71.061803 ; PhysRevA.75.062325 ; Cheng2023 ; Lita:08 ; Calkins:13 ; Reddy:20 ; Marsili2013 ; 10.1063/1.5000001 ; Zhang2017 ; Pernice2012 ; doi:10.1021/acsphotonics.4c01680 ; 10.1063/1.1388868 in recent years that can be used to perform these tests.
As another contribution, also from the manufacturer’s perspective, we derive the photon statistics for two-photon arrival-time-based QRNGs in practical scenarios. For the practical scenario, we consider power instability as the error of the source and the attenuators and detection inefficiency as the error of a single photon detector (SPD).
@@
  • If the expected number of photons within a unit of time is small, the min-entropy from the quantum source would be high. However, it would increase the cost of the QRNG, and it would also reduce the random number generation rate.
  • Although a small external reference cycle of the QRNG increases the min-entropy from the quantum source, to minimize the timing error, it should be higher than a specific value. This reference cycle does not have much effect on the cost or random number generation rate of the QRNG.
@@
Finally, we simulated raw data depending on photon statistics we have derived and performed statistical tests like NIST 10.5555/2206233 , Dieharder Die , AIS-$31$ killmann2011proposal and ENT ENT on those raw data. We show that these statistical tests support our theoretical deduction regarding randomness from quantum sources. Nie et al. 10.1063/1.4863224 performed NIST tests after post-processing and showed that the post-processed data passed all of the 15 tests from the NIST suite.
@@
This paper is structured as follows. In Section II we discuss the quantumness of a single-photon source. In Section III and IV we explore single-photon-based QRNGs and coherent pulse-based QRNGs respectively and mention their weakness in present days. We also discuss noisy beam-splitters and detectors in the case of single-photon-based QRNG in Section III. Section V reveals our two-fold proposal to check the quantumness of a photon-based quantum device as an answer to Q1. Detailed discussion on coherent pulse-based QRNGs has been performed in Section VI to answer Q2. In this section, we have also mentioned how our findings are different from the work of Nie et al. 10.1063/1.4863224 . Section VII answers Q3 by separately analyzing the effects of the parameters mentioned in that question. The answer to Q3 by combining the randomness with random number generation rate and entropy for maximum rate is presented in Section VIII. Results from different randomness testing suites for our simulated random numbers based on the analysis of Section VI are given in Section IX. Finally, in Section X we conclude our work.
II Quantumness of a Single-photon Source
In this section, we describe what we mean by quantumness in the case of a single-photon source. We first discuss photon statistics and then revisit the existing procedure to distinguish quantum light from classical light using such photon statistics. Here, we validate the quantumness of a photon source by its statistical closeness to an ideal single-photon source. Another way of defining and validating quantumness is through Bell test, quantum discord, non-separability, etc. acin2016certified ; Alicki_2008 ; Alicki_2008a ; Martin_2023 for QRNGs that are based on entangled photon source. As our point of interest is the single-photon source which does not have any entanglement, we are not concerned with these tests in our current work.
Photon detection and counting is an important part of QRNG which is further used in other practical applications such as quantum cryptography and quantum communication protocols Bennett ; Crypto . In general, photon-based QRNGs generate random numbers by detecting single photons. Therefore, photon-counting statistics play an important role in determining whether or not light falls in the quantum regime.
The distribution of the single-photon-counting statistics can be divided into three classes fox2006quantum :
  1. (i)
  2. (ii)
  3. (iii)
Classical wave theory of light follows Poissonian and super-Poissonian statistics fox2006quantum . On the other hand, the sub-Poissonian nature of a photon source cannot be explained by the classical wave theory, making it a clear indication of its quantum nature RevModPhys.68.127 ; qute.202100062 . A detailed discussion is given in B.
Let us consider a probability distribution given by,
$P(n)=\begin{cases}\frac{e^{-\alpha\mu}(\alpha\mu)^{\alpha n}}{(\alpha n)!},&\text{if }\alpha n\text{ is integer,}\\ 0,&\text{otherwise}\end{cases}$ (1)
for $n=0,1,2,\dots$ and some positive real numbers $\mu,\alpha$. Then the mean would be $\mu$ and variance would be $\frac{\mu}{\alpha}$. Therefore, this denotes a sub-Poissonian, Poissonian, or super-Poissonian distribution for $\alpha>1,\alpha=1,\alpha<1$ respectively.
Thermal light follows Bose–Einstein distribution fox2006quantum , given by
$P(n)=\frac{1}{\mu+1}\left(\frac{\mu}{\mu+1}\right)^{n}.$ (2)
$P(n,T)=\frac{e^{-\mu T}(\mu T)^{n}}{n!},$ (4)
Antibunching property can also be used to distinguish quantum light from classical light. Detailed discussion about this is included in B.
III QRNG based on Single-photon Source
An ideal QRNG, based on single photons, uses $/50$ 50 beam splitter jennewein2000fast , where an $R/T$ beam-splitter denotes an optical device that splits an incoming light beam into $R\%$ reflected and $T\%$ transmitted light beams. A standard technique is to place a $/50$ 50 Beam-splitter making an angle of $45^{\circ}$ with the path of the photons. Thus, a single photon either passes through the beam-splitter, or it is reflected with probability $\frac{1}{2}$. Two photon detectors say $D_{0}$ and $D_{1}$, can be placed over two output paths. If a photon is detected by $D_{i}$, $i$ is registered as a random bit. A block diagram of this process is given in Fig. 1.
[…]
Figure 1: A generic block diagram of the single-photon-based QRNG. Ideally, single-photon, coming from the source, are sent through a $/50$ 50 beam-splitter. Two SPDs ($D_{0},D_{1}$) are placed in two paths to detect the photon. Photon may take either path with probability $0.5$. The random bit would be $i$ if the detection happens at $D_{i}$. In practice, beam splitters and detectors cannot be error-free. However, for an $R/T$ beam-splitter, by choosing detectors as in Eq. (7), uniform random numbers can be generated.
III.1 Effect of Imperfect Devices
If the photon source for this QRNG is not a perfect single-photon source, it may happen that at the same time some photons are transmitted through the beam-splitter and some are reflected. Then, both the detectors may detect photons simultaneously, and the random number cannot be generated. To ensure whether the source is a single-photon source or not, one may perform the first step of the two-fold statistical method mentioned in Section V. If this method reveals that the distribution is not sub-Poissonian, then the source cannot be a single-photon source, and that source cannot be used in this QRNG.
Each bit of the random number generated with the above single-photon-based QRNG has probability $\frac{1}{2}$ of being 00 or $1$ when all the underlying hardware components such as source, beam-splitter, SPD, are ideal. So, ideally, it produces uniform random numbers. However, any defect in the beam-splitter or the detectors may change this probability. Let us consider an $R/T$ beam-splitter and the detectors $D_{i}$ with detection efficiency $d_{i}\in(0,1]$. Suppose, $D_{0}$ is placed on the path of the transmitted light, and $D_{1}$ is placed on the path of the reflected light. Then the probability of 00 in the generated random number, that is, the probability of detecting a photon by detector $D_{0}$ would be
$P(D_{0})=\frac{d_{0}}{d_{0}+d_{1}}\frac{T}{100}.$ (5)
Similarly, the probability of $1$ would be
$P(D_{1})=\frac{d_{1}}{d_{0}+d_{1}}\frac{R}{100}.$ (6)
Therefore, to generate random numbers, we require,
$P(D_{0})=P(D_{1})\implies d_{0}T=d_{1}R.$ (7)
If $d_{0}T\neq d_{1}R$, the generated random number would not be uniformly random. To test this randomness, statistical tests like NIST, Dieharder, AIS-$31$, ENT etc. may be used. See Section IX for a discussion on these.
IV Existing QRNG Models based on Coherent Pulse

### v6

@@ Abstract
From the user’s perspective, the fundamental parameters of a QRNG are quantumness, efficiency (entropy and random number generation rate), and cost. Our analysis reveals that these parameters depend on three factors, namely, expected photon count per unit time, external reference cycle duration, and detection efficiency. A lower expected photon count enhances entropy but increases cost and decreases the generation rate. A shorter external reference cycle boosts entropy but must exceed a minimum threshold to minimize timing errors, with minor impacts on cost and rate. Lower detection efficiency enhances entropy and lowers cost but reduces the generation rate. Finally, to validate our results, we perform statistical tests like NIST, Dieharder, AIS-31, ENT etc. over the data simulated with different values of the above parameters. Our findings can empower manufacturers to customize QRNGs to meet user needs effectively.
Keywords: Quantumness, Randomness, QRNG, Photon statistics, Entropy, Rate
@@ I Introduction
Random numbers play a vital role in various applications, including Monte Carlo simulations, software development, physics and chemistry experiments, cryptography, and commercial uses like slot machines and lotteries RevModPhys.87.1067 ; MonteCarlo ; PhysicsPhysiqueFizika.1.195 ; Bennett ; Crypto ; konstantinou2004electronic ; yamanaka2017enumeration . In cryptography, secure communication relies on high-quality randomness. Classical pseudo-random number generators (PRNGs), being deterministic, lack true unpredictability. In contrast, true random number generators (TRNGs) derive randomness from physical entropy sources like quantum effect, noise, free oscillations, chaos, etc. Schindler2009 ; Stipčević2014 ; MA2024115310 ; ZHAO2023113407 ; Ma_2025 .
Classical TRNGs rely on the unpredictability of physical processes due to incomplete knowledge Stipčević2014 ; mannalatha2023comprehensive , but future research may render these processes predictable. Quantum TRNGs (QRNGs), however, leverage quantum mechanics to generate fundamentally unpredictable numbers ma2016quantum ; Zhang2023 , ensuring lasting randomness.
@@
Different technologies, including photon arrival time 10.1063/1.2720728 ; 10.1063/1.2961000 ; WayneRNG ; WayneQRNG ; 10.1063/1.3578456 ; 10.1063/1.4863224 ; 7270265 ; zhang2021simple ; solymos2024correlation , tunneling bernardo2017extracting ; 10.1063/5.0055955 , photon detection acsomega ; wang2025efficient , fluctuation gabriel2010generator ; PhysRevApplied.3.054004 ; Guo:19 ; ZQVR ; 10.1063/1.5078547 ; gehring2021homodyne , phase noise Qi:10 ; PhysRevE.81.051137 ; PhysRevA.81.063814 ; Xu:12 ; 10.1063/1.4922417 ; Yang:16 ; app10072431 , etc. have been proposed to produce QRNGs RevModPhys.89.015004 ; mannalatha2023comprehensive . Several manufacturers such as ID Quantique quantis , Qutools quRNG , QNuLabs TROPOS , QuintessenceLabs qStream , Mars Innovation quantoss , Quantum Computing Inc. uQRNG , Quside quside , QuantumCTek QCTek are producing QRNGs following these technologies. These QRNGs are commercially available as off-the-shelf black-box devices. As the photon arrival-time-based QRNGs are the most general and most of companies follow this approach, in this work we focus on the analysis of this type of QRNGs only.
@@
Since no black-box randomness-testing algorithm can distinguish between a pseudorandom number generator (PRNG) and a quantum random number generator (QRNG) PhysRevA.109.022243 , the first step in validating the quantumness of a random number generator is to gather detailed information about the intermediate components, such as attenuator and detector parameters, and photon statistics. This data may then enable a strategy to determine whether the device operates in the classical or quantum regime. Here, by quantumness we denote whether the underlying photon-counting-statistics follows sub-Poissonian distribution ($\text{mean}>\text{variance}$). For example, for a probability distribution given by $P(n)=\frac{e^{-\alpha\mu}(\alpha\mu)^{\alpha n}}{(\alpha n)!},\;\text{if }\alpha n\text{ is integer, and }0$ otherwise, for $n=0,1,2,\dots$, where $\mu,\alpha$ are some positive real numbers, the mean would be $\mu$ and the variance would be $\frac{\mu}{\alpha}$. Thus it denotes a sub-Poissonian, Poissonian, or super-Poissonian distribution for $\alpha>1,\alpha=1,\alpha<1$, respectively. Therefore, if $\alpha>1$, we can surely say that this falls under the quantum regime. We provide real-world examples in A.
Numerous photon-number-resolving detectors have been proposed PhysRevA.71.033819 ; Divochiy2008 ; Kardyna2008 ; PhysRevA.99.043822 ; Natarajan2012 ; Schuck2013 ; Hadfield2009 ; 9844820 ; Provaznik:20 ; Marsili2009 and experimentally demonstrated 1263782 ; PhysRevA.67.061801 ; 10.1063/1.1596723 ; PhysRevA.71.061803 ; PhysRevA.75.062325 ; Cheng2023 ; Lita:08 ; Calkins:13 ; Reddy:20 ; Marsili2013 ; 10.1063/1.5000001 ; Zhang2017 ; Pernice2012 ; doi:10.1021/acsphotonics.4c01680 ; 10.1063/1.1388868 in recent years, providing photon-counting statistics for analysis of the underlying distribution. In most analyses of photon arrival-time-based QRNGs Jones:23 ; Perina:24 ; Perina24 ; PhysRevA.109.043505 , photon statistics are examined by comparing the sample mean and variance or using the Mandel parameter Mandel:79 . For instance, Alléaume et al. Alléaume_2004 analyzed photon count data using the Mandel parameter (Mandel:79, , Eq. (11a)), concluding that its negative value, which indicates sub-Poissonian statistics, signifies non-classical photon sources. While analytically valid, this approach may be misleading for finite samples. A negative Mandel parameter can appear even for Poissonian sources like coherent light due to finite-size sampling error, where the sample mean and variance deviate from those of the true distribution BERG2005865 . Thus, a single parameter value from a finite dataset is insufficient to infer the underlying photon statistics.
Another method to assess the quantumness of light is through photon antibunching RevModPhys.54.1061 . Although antibunched photons indicate non-classicality, the absence of antibunching does not imply classical behavior; non-classical light may exhibit photon bunching PhysRevA.41.475 ; fox2006quantum . This criterion also relies on statistical equality, making it similarly inconclusive in certain cases. A detailed discussion appears in Section A.
On the other hand, a single-photon source ensures the quantumness of a QRNG RevModPhys.54.1061 ; fox2006quantum , though such sources remain experimentally challenging PhysRevLett.72.210 ; kim1999single . Various approaches have been proposed to generate single photons PhysRevLett.83.2722 ; lounis2000single ; PhysRevLett.86.1502 ; santori2002indistinguishable ; 10.1063/1.1577828 ; PhysRevLett.89.067901 ; adam2024single ; wang2024multimode .
Time-bin-based QRNGs typically use attenuated coherent sources to approximate single-photon emission and generate random numbers with exponential WayneQRNG ; WayneRNG ; 10.1063/1.3578456 or uniform 10.1063/1.4863224 distributions under ideal conditions. However, in practice, deviations occur. For example, Nie et al. 10.1063/1.4863224 demonstrated that uniformity, and thus min-entropy, depends on parameters like photon numbers and detection efficiency. This highlights the importance of analyzing photon statistics under realistic conditions.
Although QRNGs aim to produce uniformly distributed outputs, imperfections in hardware introduce deviations. As a result, classical post-processing is required to extract uniform randomness. It is important to note that such processing does not increase entropy, but compresses the raw output to extract existing min-entropy, often at the cost of reduced generation rates.
@@ I.1 Motivation
Several device-independent QRNG (DI-QRNG) schemes pironio2010random ; 10.1145/2885493 ; acin2016certified ; bierhorst2018experimentally ; liu2018device ; PhysRevLett.120.010503 ; 9153851 and certified randomness protocols pironio2010random ; 6326426 ; Passaro_2015 ; acin2016certified ; Nguyen:18 ; 10.1093/oso/9780198788416.003.0008 ; PhysRevX.10.041048 ; zhang2021simple ; Leone:23 ; PhysRevLett.133.020802 ; PRXQuantum.5.020348 based on Bell violation and non-separability acin2016certified ; Alicki_2008 ; Alicki_2008a ; Martin_2023 have been proposed. These approaches do not rely on trusted devices or require post-processing, but they demand costly quantum resources such as entangled qubits. Recent efforts Mironowicz_2016 ; 10.1145/3564246.3585145 explore locally certified randomness without entanglement, though they still require random seeds and bootstrapping steps, functionally similar to classical post-processing.
In contrast, optical trusted-device QRNGs mannalatha2023comprehensive assume that both the source and detection setup are well-characterized and isolated from external influence. While less robust than DI-QRNGs, they are more practical and can be verified using photon-count statistics PhysRevA.41.475 ; fox2006quantum ; RevModPhys.68.127 ; qute.202100062 , offering enhanced security under trusted assumptions. A detailed comparison is provided in Section B. This work focuses on optical trusted-device QRNGs, and introduces a two-fold statistical approach to verify their quantumness.
Side-channel attacks 8924447 ; 8942152 ; chowdhury2021physical may target the post-processing stage to extract information from the output. Hence, minimizing post-processing is essential to reduce compression overhead, improve throughput, and mitigate attack surfaces. For this, the raw output should be as close to uniform as possible. Nie et al. 10.1063/1.4863224 derived a lower bound on the min-entropy of raw output under device imperfections. Based on this min-entropy, suitable post-processing can be applied to retain entropy from the quantum source. As all practical systems involve errors, this raises a critical question:
@@ I.2 Our Contributions
In this paper, we aim to evaluate the quantumness of photon arrival-time-based QRNG from the manufacturer’s perspective as answers to Q1 and Q2 and randomness, random number generation rate, and cost from the user’s perspective as the answer to Q3. A similar analysis can also be performed for other QRNG models. For the manufacturer, we first address the validation of quantum sources using photon counting statistics. Considering that the sample data for the number of photons within a fixed time is available, we suggest a two-fold statistical method to verify whether the source of the sample is quantum or not, overcoming the limitations of the existing approaches discussed above. The method consists of interval estimation and hypothesis testing to ensure sub- or super-Poissonian distribution up to a desired confidence level. Several photon-number-resolving detectors have been proposed PhysRevA.71.033819 ; Divochiy2008 ; Kardyna2008 ; PhysRevA.99.043822 ; Natarajan2012 ; Schuck2013 ; Hadfield2009 ; 9844820 ; Provaznik:20 ; Marsili2009 and demonstrated 1263782 ; PhysRevA.67.061801 ; 10.1063/1.1596723 ; PhysRevA.71.061803 ; PhysRevA.75.062325 ; Cheng2023 ; Lita:08 ; Calkins:13 ; Reddy:20 ; Marsili2013 ; 10.1063/1.5000001 ; Zhang2017 ; Pernice2012 ; doi:10.1021/acsphotonics.4c01680 ; 10.1063/1.1388868 in recent years that can be used to perform these tests.
As another contribution, also from the manufacturer’s perspective, we derive the photon statistics for two photon arrival-time-based QRNGs in practical scenarios. We consider power instability as the error of the source and the attenuators and detection inefficiency as the error of a single photon detector (SPD).
@@
  • If the expected number of photons within a unit time is small, the min-entropy from the quantum source would be high. However, it would increase the cost of the QRNG, and it would also reduce the random number generation rate.
  • Although a small external reference cycle of the QRNG increases the min-entropy from the quantum source, to minimize the timing error, it should be bounded by a specific value. This reference cycle does not have much effect on the cost or random number generation rate of the QRNG.
@@
Finally, we simulated raw data depending on the photon statistics we have derived and performed statistical tests like NIST 10.5555/2206233 , Dieharder Die , AIS-$31$ killmann2011proposal and ENT ENT on those raw data. We show that these statistical tests support our theoretical deduction regarding randomness from quantum sources. Nie et al. 10.1063/1.4863224 performed NIST tests after post-processing and showed that the post-processed data passed all of the 15 tests from the NIST suite.
@@
This paper is structured as follows. In Section II we explore coherent pulse-based QRNGs and mention their weakness. Section III reveals our two-fold proposal to check the quantumness of a photon-based quantum device as an answer to Q1. Detailed discussion on coherent pulse-based QRNGs has been performed in Section IV to answer Q2. In this section, we have also mentioned how our findings are different from the work of Nie et al. 10.1063/1.4863224 . Section V answers Q3 by separately analyzing the effects of the parameters mentioned in that question. The answer to Q3 by combining the randomness with random number generation rate and entropy for maximum rate is presented in Section VI. Results from different randomness testing suites for our simulated random numbers based on the analysis of Section IV are given in Section VII. Finally, in Section VIII we conclude our work.
II Existing QRNG Models based on Coherent Pulse